# Activation degree thresholds and expressiveness of polynomial neural networks

### v3

@@ 4 Outlook
Our bound in Theorem 18 on the activation degree threshold for polynomial neural networks is not universally optimal. We proved this to be the case for equi-width networks in Theorem 22. One way to improve the universal bound of activation degree thresholds is to leverage other algebro-geometric and number theoretic results. For instance, we expect that a lemma of [17] used to prove Picard theorems for holomorphic maps into Fermat hypersurfaces will allow us to improve the bound in Theorem 18.
@@
One future direction in this research area is to determine the activation degree thresholds for non-increasing architectures $\mathbf{d}=(d_{0},d_{1},\dotsc,d_{L})$ — those architectures satisfying $d_{0}\geq d_{1}\geq\dotsb\geq d_{L}$. By generalizing the proof of Theorem 22, one might hope to prove [28, Conjecture 5.7] that for a non-increasing architecture $\mathbf{d}$ with $d_{L}>1$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension for all $r>1$. That is, prove that a non-increasing architecture $\mathbf{d}$ has activation degree threshold less than or equal to one.
@@
Another research direction is to introduce the notion of activation degree thresholds for other linear and non-linear neural networks. Permitting negative exponents in a polynomial activation map leads to the notion of rational neural networks [8]. There is experimental evidence that it is beneficial to choose rational activation maps that are very low degree. Defining activation degree thresholds for rational networks would provide a quantitative measure to make these observations rigorous. Rational neural networks are currently being investigated by Alexandros Grosdos et al. Certain linear networks may also admit the definition of an activation degree threshold as a means to study expressiveness. Tropical rational neural networks provide a lens to study networks with piecewise linear activation maps [3, 11, 12, 45]. Such networks encompass ReLU networks; a ReLU network of fixed architecture is described by a semialgebraic set inside the parameter space of tropical rational functions. In this context, tropical rational functions provide a means to define activation degree thresholds to frame and obtain results on expressiveness. Moreover, because the dimension of the neurovariety is preserved under tropicalization, activation degree thresholds may aid in translating expressiveness results for tropical neural networks to polynomial networks.
@@ Acknowledgments
The authors would like to thank Jiayi Li, Kaie Kubjas, and Alexandros Grosdos for their comments and feedback that led to improvements of this article.

### v4

@@ 4 Outlook
Our bound in Theorem 18 on the activation degree threshold for polynomial neural networks is not universally optimal. We proved this to be the case for equi-width networks in Theorem 22. Improving the bound established in Lemma 15 leads to better bounds on the activation degree threshold in Theorem 18. One potential way to improve the universal bound of activation degree thresholds is to leverage other algebro-geometric and number theoretic results. Specifically, if the polynomials $p_{1},\dotsc,p_{k}$ in Lemma 15 satisfy the additional hypothesis of having no common zeros, then the result in [16, page 71] implies the improved bound $r\leq k^{2}-1$. In the language of algebraic geometry, [16, page 71] establishes a Picard theorem for holomorphic maps into Fermat hypersurfaces. We are hopeful that other results in this spirit will be applicable to the study of the dimension of neurovarieties.
@@
One future direction in this research area is to determine the activation degree thresholds for non-increasing architectures $\mathbf{d}=(d_{0},d_{1},\dotsc,d_{L})$ — those architectures satisfying $d_{0}\geq d_{1}\geq\dotsb\geq d_{L}$. By generalizing the proof of Theorem 22, one might hope to prove [27, Conjecture 5.7] that for a non-increasing architecture $\mathbf{d}$ with $d_{L}>1$, the neurovariety $\mathcal{V}_{\mathbf{d},r}$ has the expected dimension for all $r>1$. That is, prove that a non-increasing architecture $\mathbf{d}$ has activation degree threshold less than or equal to one.
@@
Another research direction is to introduce the notion of activation degree thresholds for other neural networks. Permitting negative exponents in a polynomial activation map leads to the notion of rational neural networks [8]. There is experimental evidence that it is beneficial to choose rational activation maps that are of very low degree. Defining activation degree thresholds for rational networks would provide a quantitative measure to make these observations rigorous. Rational neural networks are currently being investigated by Alexandros Grosdos et al. Certain piecewise linear networks may also admit the definition of an activation degree threshold as a means to study expressiveness. Tropical rational neural networks provide a lens to study networks with piecewise linear activation maps [3, 10, 11, 46]. Such networks encompass ReLU networks; a ReLU network of fixed architecture is described by a semialgebraic set inside the parameter space of tropical rational functions. In this context, tropical rational functions provide means to define activation degree thresholds to frame and obtain results on expressiveness. Moreover, because the dimension of the neurovariety is preserved under tropicalization, activation degree thresholds may aid in translating expressiveness results for tropical neural networks to polynomial networks.
@@ Acknowledgments
The authors would like to thank Jiayi Li, Kaie Kubjas, and Alexandros Grosdos for their comments and feedback that led to improvements of this article. We also thank the referees for their thoughtful suggestions and questions.